# Compressing Transformer-Based Self-Supervised Models for Speech Processing

### v2

@@ IV Design of Experiments
Compressing self-supervised models inherits the same goal of self-supervised learning, i.e., to provide representations for downstream tasks. To measure the trade-off between compression and downstream performance, we pre-train and compress 12-layer Transformers with the entire LibriSpeech, a data set consisting of 960 hours of read speech. For the downstream tasks, we conduct phone recognition on the 100-hour subset of LibriSpeech, and speaker identification on VoxCeleb 1. All the experiments follow default SUPERB settings [1], where the pre-trained models are frozen after pre-training and are not fine-tuned. Before making predictions, a downstream classifier learns a weighted sum of all 12 layers in a Transformer. The weights are learned together with the downstream models.
When measuring the amount of compression, we define density as the amount pruned relative to the total amount that can be pruned. For example, the density in head pruning is the amount of heads pruned over the total number of heads.
In this section, we will detail the training hyperparameters and show a few results that shape our design decisions. A more comprehensive comparison among all approaches will follow in the next section.
@@ IV-A MelHuBERT
We train MelHuBERT-10ms and MelHuBERT-20ms with masked prediction [3, 4] for two pre-training stages. For the first stage, we first run k-means with 512 clusters on log Mel features extracted from the 960-hour subset of LibriSpeech. We then train a 12-layer Transformer to predict the cluster labels of each frame for MelHuBERT-10ms or to predict every other frame for MelHuBERT-20ms. For the second stage, we run k-means with 512 clusters on the ninth layer hidden representations of the first stage model. We predict each frame for both MelHuBERT-10ms and MelHuBERT-20ms in the second stage. We choose a masking strategy based on [28], using a 7% masking probability with a mask of 10 frames (100 ms) for MelHuBERT-10ms and a 14% masking probability with a mask of 5 frames (100 ms) for MelHuBERT-20ms. To compare, HuBERT uses 8% of masking probability with a mask of 10 frames (200 ms) [3]. For the first stage, both MelHuBERT-10ms and MelHuBERT-20ms are trained on 8 V100 GPUs with Adam for 200 epochs. For the second stage, we train MelHuBERT-10ms for 200 epochs and MelHuBERT-20ms for 100 epochs. We use an effective batch size of 32 and a learning rate of $10^{-4}$ for both stages. It requires about 32 days to train MelHuBERT-10ms in total, and about 15 days for MelHuBERT-20ms. Dropout of 10% is applied after the multiplication of matrices, such as query, key, and value matrices, and after FC1 and FC2.
@@ IV-B Weight Pruning
Initial results of pruning MelHuBERT-10ms are shown in Figure 1. As the weights are pruned (reading from high density to low density), the pre-training loss first improves and later degrades. The improvement of the loss suggests that lightly pruning a pre-trained model has a regularization effect, and this finding will come up again in other experiments. We also find that the loss is indicative of the downstream performance. The density when the loss starts to degrade too much is also the point where the downstream performance starts to degrade. These results, though intuitive, are largely missing in prior work in the context of pruning self-supervised models with the pre-training loss.
Pre-training lossPhone recognitionSpeaker identification
Figure 1: Losses and downstream performance of weight pruning on MelHuBERT-10ms at different model densities. The dashed lines are the results of the unpruned MelHuBERT-10ms (i.e., with a 100% density).
@@ IV-C Head Pruning
For head pruning, recall that we have two approaches, one based on the $\ell_{1}$ norm of head weights, and the other based on the $\ell_{1}$ norm of gradients. For the weight-based approach, we prune a fixed amount of heads for each layer, because higher layers tend to have larger $\ell_{1}$ norm. If we compare heads across layers, an entire layer would be pruned before we prune others. For the gradient-based approach, we use a 25% of the training data to compute the scores of each head. Similarly, since the $\ell_{1}$ norm varies across layers, as opposed to pruning a fixed number for each layer, we normalize the scores of heads within each layer and prune the heads by comparing them altogether. As for all iterative pruning approaches, we train the model for a fixed 25,000 steps after pruning, with learning rate $10^{-5}$ and a batch size of 4.
Performance on downstream tasks behave similarly to weight pruning, and we defer the comparison until the next section. Results of comparing the two approaches of head pruning are shown in Figure 2. We find that the gradient-based approach is better than the weight-based approach, and will focus on the gradient-based approach for the rest of the paper.
weight-based head pruninggradient-based head pruning
Refer to caption Refer to caption
(a) phoneme recognition (b) speaker identification
Figure 2: Phoneme recognition and speaker identification performance comparing weight-based and gradient-based head pruning on MelHuBERT-10ms. The dashed line is the performance of the unpruned MelHuBERT-10ms.
IV-D Low-Rank Approximation
For low-rank approximation, recall that we target the 3072-dimensional output of the FC1 layer. Reducing the dimension amounts to pruning the rows of FC1 and the columns of FC2 based on the sum of their $\ell_{1}$ norm. The input and output dimensions (in our case 768) are seldom larger than the dimension in the middle (in this case 3072).
Before any pruning, we first inspect whether these matrices are low-rank after pre-training. Figure 3 shows the singular values of FC1 and FC2. From the sharp drop in values after the first few dimensions, it is clear that nearly all matrices in FC1 and FC2 are low-rank. We would expect that a lot of the rows and columns can be pruned. We prune 128 dimensions every 25,000 steps, and use a learning rate of $10^{-5}$ and a batch size of 4 to train the model after every pruning step. Performance on downstream tasks and comparison to other approaches are deferred to the next section.
Refer to caption Refer to caption
(a) Singular values of FC1 (b) Singular values of FC2
Figure 3: Singular values of the FC1 and FC2 weight matrices in MelHuBERT-10ms. In each figure, the x-axis is the number of dimension (in this case 768) and the y-axis is the value itself.
IV-E Knowledge Distillation
For knowledge distillation, we explore a 2-layer and a 6-layer student network. Prior work [20, 26] has suggested that some masking of the input would be beneficial for distillation as opposed to not masking. The same study also suggest initializing student networks with middle layers taken directly from teacher networks. The temperature of the teacher distribution is also encouraged to be higher than 1, because the output distribution tends to be peaky [12]. Setting the temperature high would make the values of the tail larger. Given these reasons, we study several design decisions, including changing the masking strategy, initializing with MelHuBERT layers, and the temperature when training with KL divergence. Based on our study, neither applying masking on the input nor initializing with MelHuBERT layers lead to improvement. Hence, we simply minimize KL divergence as our distillation loss without masking and initializing.
Results of compressing MelHuBERT for both 10 ms and 20 ms frame periods are shown in Figure 4. We find that both the 2-layer and 6-layer students are able to retain the performance of teachers on speaker identification, with the 6-layer ones performing even better than the teachers. Both students are not able to match the teachers’ performance on phone recognition, with the 2-layer one lacking significantly behind. The worse performance seems to suggest that knowledge distillation might not be an effective compression approach. However, as we will see in later sections, once we put all the results in context, knowledge distillation actually performs favorably compared to others.
@@ IV-E Knowledge Distillation
V Main Results
@@ V-A Downstream performance
We first study downstream performance, specifically phone recognition and speaker identification, along the pruning process, i.e., at different densities. Note that density is measured relative to the total amount that can be pruned. For example, in head pruning, density is the number of attention heads left over the total number of heads in the unpruned Transformer. We are interested in to what extent downstream performance is maintained while we prune models with the pre-training loss. We compare approaches that belong to the iterative pruning family, namely, weight pruning, head pruning, and low-rank approximation. Knowledge distillation will be compared in the next section. Results are shown in Figure 5.
@@
V-B MACs, parameterss, and runtime
While Figure 5 shows how well each approach maintains performance along the pruning process, it does not show the benefit of compression. Density is a relative measure, making it difficult to compare across approaches. To evaluate the benefit of compression across approaches, we decide to use three metrics, the number of multiply-accumulation operations (MACs) per one second speech, the number of parameters, and the runtime measured on a single GPU. The three metrics are related. Reduction in MACs represents the theoretical speed-up of inference when parallel computation is not available. In the presence of a GPU, the reduction in MACs does not always translate to faster runtime, as it depends on the structure of the pruned networks. Reduction in model parameters usually leads to a reduction in MACs and can improve generalization. As we have seen in weight pruning and head pruning (Figure 5), there is a period along the pruning process where the pre-trained loss and downstream performance both improve. Results of all approaches evaluated on the three metrics are shown in Figure 6. We only focus on phone recognition, but it shows a similar trend for speaker identification.
@@ V-B MACs, parameterss, and runtime
Knowledge distillation (with a 2-layer and a 6-layer student) achieves a significant reduction on all three metrics. In fact, knowledge distillation provides the best trade-off among all approaches. We attribute this to the the number of GPU kernel calls. Each GPU kernel call has a significant overhead, and computing with fewer layers not only reduces MACs per one second speech, but also saves the overhead. For example, a model pruned with low-rank approximation can have a similar MACs per one second speech compared to a 6-layer student, but the runtime of the pruned model is significantly slower than the 6-layer student.
@@
Finally, inspired by the results of knowledge distillation, we provide a baseline that only takes the first two layers and the first six layers of the pre-trained Transformers. This amounts to stopping the forward process at the second and the sixth layer. As we have argued, stopping the forward process can be seen as a form of compression. This approach involves no additional training and is arguably the simplest baseline. Combined with other approaches in Figure 6, we now have a better picture what the worst and the best are able to achieve. These results enable a wide range of applications, and we will explore a few in subsequent sections.
VI Applications
In the previous section, we have shown the trade-off of four model compression approaches, namely, weight pruning, head pruning, low-rank approximation, and knowledge distillation on three different metrics, MACs per one second speech, numbers of parameters, and real-time factors. In this section, we will showcase a few applications of our findings, including contextualizing prior work, designing new compression techniques, and analyzing Transformers.
VI-A Contextualizing prior work
There have been several attempts compressing self-supervised Transformers, such as wav2vec 2.0 and HuBERT. However, most studies focus on the breadth of tasks that can be maintained after compression, rather than the trade-off discussed in the previous section. To fill in this gap, we choose three approaches based on compressing HuBERT, i.e., DistilHuBERT [19], FitHuBERT [21], and LightHuBERT [20]. Given our findings, we can not only discuss their strengths and weaknesses, but also explain why.
Figure 7 shows how DistilHuBERT, FitHuBERT, and LightHuBERT (small) compare against our compression approaches on phone recognition. First, we note that LightHuBERT has the best PER. In fact, LightHuBERT has better PER before compressing, due to an initial round of self-distillation.
DistilHuBERT and LightHuBERT both consumes significantly more MACs per one second speech, as they both retain the convolutional layers from HuBERT. In LightHuBERT for example, the convolutional layers accounts for about 64% of MACs while only constituting 4.4% of the parameters. On the other hand, the convolutional layers in FitHuBERT are optimized; hence consuming significantly fewer MACs per one second speech.
@@ VI-A Contextualizing prior work
The three have similar numbers of parameters but significantly different performance on phone recognition. Since DistilHuBERT has 2 layers while FitHuBERT and LightHuBERT have 12, the result suggests that, for phone recognition, the depth of Transformers is more critical than the number of heads and the dimensions in the feed-forward layers.
In terms of the actual runtime however, LightHuBERT is the worst while FitHuBERT is not better than simply taking the first six layers of MelHuBERT. As we have argued in the previous section, using more layers leads to more GPU kernel calls. The overhead of GPU kernel calls dominates especially when each kernel is cheap to compute.
In sum, our results provide context for DistilHuBERT, FitHuBERT, and LightHuBERT, and can explain their stengths and weaknesses. In addition, simple baseines are strong compared to these recent approaches.
VI-B Combining compression techniques
@@
VI-C Analyzing weights pruned
Model compression not only is an approach for reducing model size, but also an approach to understand the internals of Transformers. In this section, we analyze where the weights are pruned during weight pruning. Given that different layers have different downstream performance [13] or certain Transformer heads [32] are specialized, it is likely that weight pruning would help us identify parts of Transformer that are redundant or not useful for maintaining the pre-training loss. First, Figure 8 (a) shows the density of layers along the process of weight pruning. To our surprise, the layers are equally pruned along the process of weight pruning, with the last layer pruned slightly more compared to others. We then measure the density for different weight matrices, including the matrices for computing queries, keys, and values, and the matrices in FC1 and FC2. As shown in Figure 8 (b), again, to our surprise, weight matrices are mostly equally pruned throughout the pruning process, with the matrix for computing the values slightly more pruned than others.
Our results suggest several new research directions. Recall that we can maintain the pre-training loss with weight pruning until at least a density of 40%. Weight pruning is able to find a Transformer that is sparse uniformly compared to the unpruned one. It would be interesting to know whether the uniformity is due to the pruning algorithm. Given the strong performance of weight pruning, it’s also interesting to know whether the uniformity is necessary or even optimal for pruning.
[…]
Refer to caption Refer to caption
(a) Layers (b) Matricies
Figure 8: Left: The average densities of layers during weight pruning. The lighter the color, the lower the layer. Right: The average densities of specific weight matrices during weight pruning.
Figure 9: Left: The average diagonality of Transformer heads at different densities during head pruning with the gradient-based approach. Center and Right: Phone recognition results of head pruning by approximating highly diagonal attention maps with identity matrices and 1-D convolutions. Gradient-based head pruning is also shown for reference.
VI-D Analyzing heads pruned
In addition to analyzing weight pruning, we similarly use model compression to analyze head pruning. Several studies have observed that attention maps tend to be diagonal after training [51, 52, 53]. If the attention maps are exactly diagonal, then self-attention is nothing but a feedforward layer and can be replaced with a feedforward layer to save computation [51]. In fact, if there exist many diagonal attention heads, they might be redundant and potentially be pruned during head pruning.
To study what attention heads are deemed more important and its relationship to diagonality, we follow [51] and measure diagonality of attention heads in each pruning iteration. Specifically, for a position $i$, the attention weight $a_{ij}$ to position $j$ is scaled by $|i-j|$ to measure the weight that is off diagonal. The scaled attention is then normalized by the largest number of positions possible on either side of $i$, i.e., $\max(|i-1|,\dots,|i-T|)$, where $T$ is the sequence length. The diagonality is defined as
$1-\frac{1}{T}\sum_{i=1}^{T}\frac{\sum_{j=1}^{T}a_{ij}|i-j|}{\max(|i-1|,\dots,|i-T|)}.$ (9)
We calculate diagonality for each attention head and average over all attention heads. We then average the diagonality of a pruned model over 100 utterances. Results are shown in the left of Figure 9. The diagonality of the unpruned Transformer is about 0.64, a high number consistent with the findings of others. However, to our surprise, the diagonality increases when more heads are being pruned. The increase in diagonality is due to the fine-tuning inbetween pruning. Attention heads that are less diagonal becomes more diagonal after fine-tuning. This suggests that diagonal attention heads have their significance in Transformers.
Given that many attention maps become diagonal, a potential approach to compress attention heads is to replace the highly diagonal ones with identity matrices. In particular, if
$\text{softmax}\left(\frac{QK^{\top}}{\sqrt{d}}\right)\approx\mathrm{I}$ (10)
we can skip the computation of the query $Q$ and the key $K$ and directly output the value matrix $V$. Similar to regular head pruning, we iteratively replace highly diagonal attention heads with identity matrices according to diagonality in equation (9) and fine-tune the rest. We refer to this as the skip-ahead approach. The skip-ahead approach might be too crude an approximation. A weighted sum that centers around the diagonal can be approximated with a 1-D convolution. Instead of replacing a head with identity matrix, we replace it with a 1-D convolution.
Results on phone recognition for both the skip-ahead and 1-D convolution approaches are shown in the center and the right of Figure 9. The skip-ahead approach is indeed worse than regular head pruning, However, the performance is only about 2 to 3% worse, showing that the approximation does in fact align well with our findings. For the 1-D convolution, we experiment several filter sizes and find that a filter of size 21 (i.e., 10 frames on each side) performs the best. The 1-D convolution approach is on par with regular head pruning.
In conclusion, we observe that head pruning leads to an increase in diagonality during the pruning process, motivating the two approaches for pruning diagonal attention heads. Both the skip-ahead and 1-D convolution approaches successfully prune diagonal attention heads without much performance degradation. These findings could potentially inspire novel head pruning approaches.
VII Conclusion
In this paper, we explore several simple and effective approaches for pruning self-supervised Transformers, including weight pruning, head pruning, low-rank approximation, and knowledge distillation. We report MACs per one second speech, numbers of parameters, and real-time factors, and recommend others to study compression from multiple aspects, as different approaches have different strengths and weaknesses. We present several applications of our results, contextualizing recent approaches, chaining several compression techniques, discussing the redundant parts of the architecture in Transformers, and suggesting novel head pruning approaches. Model compression is not just an approach to get small and efficient models, and we hope this work will spawn more use of model compression as a tool for analyzing models.

### v3

@@ IV Design of Experiments
To understand compression tradeoffs, we pre-train and compress 12-layer Transformers using the entire LibriSpeech dataset, which contains 960 hours of read speech. We evaluate representations on two downstream tasks: phoneme recognition using the 100-hour subset of LibriSpeech, and speaker identification using VoxCeleb1. Following SUPERB [32], pre-trained models are frozen, and downstream classifiers learn a weighted sum of all layers. All pruning experiments use pre-training loss as a proxy for downstream performance, and we show loss curves and key observations to motivate our design choices.
@@ IV-A MelHuBERT
We train MelHuBERT-10ms and MelHuBERT-20ms with masked prediction [4, 33] for two pre-training stages. For the first stage, we first run k-means with 512 clusters on log Mel features extracted from the 960-hour subset of LibriSpeech. We then train a 12-layer Transformer to predict the cluster labels of each frame for MelHuBERT-10ms or to predict every other frame for MelHuBERT-20ms. For the second stage, we run k-means with 512 clusters on the ninth layer hidden representations of the first stage model. We predict each frame for both MelHuBERT-10ms and MelHuBERT-20ms in the second stage. We choose a masking strategy based on [25], using a 7% masking probability with a mask of 10 frames (100 ms) for MelHuBERT-10ms and a 14% masking probability with a mask of 5 frames (100 ms) for MelHuBERT-20ms. For the first stage, both MelHuBERT-10ms and MelHuBERT-20ms are trained on 8 V100 GPUs with Adam for 200 epochs. For the second stage, we train MelHuBERT-10ms for 200 epochs and MelHuBERT-20ms for 100 epochs. We use an effective batch size of 32 and a learning rate of $10^{-4}$ for both stages. Dropout of 10% is applied after the multiplication of matrices, such as query, key, and value matrices, and after FC1 and FC2.
@@ IV-C Head Pruning
For head pruning, we use a 25% of the training data to compute the scores of each head. Similarly, since the $\ell_{1}$ norm varies across layers, as opposed to pruning a fixed number for each layer, we normalize the scores of heads within each layer and prune the heads by comparing them altogether. As for all iterative pruning approaches, we train the model for a fixed 25,000 steps after pruning, with learning rate $10^{-5}$ and a batch size of 4.
Pre-training lossPhone recognitionSpeaker identification(a) 10 ms
Refer to caption(b) 20 ms
@@
IV-D Low-Rank Approximation
For low-rank approximation, recall that we target the 3072- dimensional output of the FC1 layer. Reducing the dimension amounts to pruning the rows of FC1 and the columns of FC2 based on the sum of their $\ell_{1}$ norm. The input and output dimensions (in our case 768) are seldom larger than the dimension in the middle (in this case 3072). We prune 128 dimensions every 25,000 steps, and use a learning rate of $10^{-5}$ and a batch size of 4 to train the model after every pruning step.
IV-E Knowledge Distillation
For knowledge distillation, we explore a 2-layer and a 6-layer student network. Following prior works [10, 11], we study several design decisions, including changing the masking strategy, initializing with MelHuBERT layers, and the temperature when training with KL divergence. Based on our study, neither applying masking on the input nor initializing with MelHuBERT layers lead to improvement. Hence, we simply minimize KL divergence as our distillation loss without masking and initializing.
V Results
@@
MelHuBERT MelHuBERT-first2 MelHuBERT-first6weight pruninghead pruninglow-rank approximationknowledge distillation
(a) 10 msRefer to caption
(b) 20 msRefer to caption
@@ V-A Downstream performance
We first study downstream performance, specifically phone recognition and speaker identification, along the pruning process, i.e., at different densities. It is noted that density is measured relative to the total amount that can be pruned. For example, in head pruning, density is the number of attention heads left over the total number of heads in the unpruned Transformer. We are interested in to what extent downstream performance is maintained while we prune models with the pre-training loss. The results of the iterative pruning family are presented in Figure 1, while those of knowledge distillation are shown in Figure 2.
@@
For knowledge distillation, we find that both the 2-layer and 6-layer students are able to retain the performance of teachers on speaker identification, with the 6-layer ones performing even better than the teachers. Both students are not able to match the teachers’ performance on phone recognition, with the 2-layer one lacking significantly behind. The worse performance seems to suggest that knowledge distillation might not be an effective compression approach. However, as we will see in later sections, once we put all the results in context, knowledge distillation actually performs favorably compared to others.
V-B MACs, parameters, and runtime
While Figure 1 and Figure 2 shows how well each approach maintains performance after compression, it does not show the benefit of compression. Density is a relative measure, making it difficult to compare across approaches. To evaluate the benefit of compression across approaches, we decide to use three metrics, the number of multiply-accumulation operations (MACs) per one second speech, the number of parameters, and the runtime measured on a single GPU. The three metrics are related. Reduction in MACs represents the theoretical speed-up of inference when parallel computation is not available. In the presence of a GPU, the reduction in MACs does not always translate to faster runtime, as it depends on the structure of the pruned networks. Reduction in model parameters usually leads to a reduction in MACs and can improve generalization. As we have seen in weight pruning and head pruning (Figure 1), there is a period along the pruning process where the pre-trained loss and downstream performance both improve. Results of all approaches evaluated on the three metrics are shown in Figure 3. We only focus on phone recognition, but it shows a similar trend for speaker identification.
@@ V-B MACs, parameters, and runtime
MelHuBERT MelHuBERT-first2 MelHuBERT-first6weight pruninghead pruninglow-rank approximationknowledge distillationknowledge distillation followed by head pruning and low-rank approximation
@@
Knowledge distillation (with a 2-layer and a 6-layer student) achieves a significant reduction on all three metrics. In fact, knowledge distillation provides the best tradeoff among all approaches. We attribute this to the the number of GPU kernel calls. Each GPU kernel call has a significant overhead, and computing with fewer layers not only reduces MACs per one second speech, but also saves the overhead. For example, a model pruned with low-rank approximation can have a similar MACs per one second speech compared to a 6-layer student, but the runtime of the pruned model is significantly slower than the 6-layer student.
@@
Finally, inspired by the results of knowledge distillation, we provide a baseline that only takes the first two layers and the first six layers of the pre-trained Transformers. This amounts to stopping the forward process at the second and the sixth layer. This approach involves no additional training and is arguably the simplest baseline. Combined with other approaches in Figure 3, we now have a better picture what the worst and the best are able to achieve.
V-C Contextualizing prior work
In addition to the four classic compression methods adopted in this study, several more advanced techniques have been proposed for compressing self-supervised Transformers. However, most of these methods emphasize maintaining performance across a wide range of tasks rather than focusing on the tradeoffs discussed in the previous section. To address this gap, we selected five HuBERT-based compression methods for comparison: DistilHuBERT [8], FitHuBERT [9], LightHuBERT [10], ARMHuBERT [11], and STaRHuBERT [12]. Our findings allow us not only to identify their respective strengths and weaknesses but also to explain the reasons behind them.
Figure 4 illustrates the performance of these five methods compared to the classic compression approaches we study, evaluated on phone recognition. Among them, LightHuBERT achieves the lowest phone error rate (PER). This is largely because LightHuBERT undergoes an initial round of self-distillation before compression, resulting in a strong baseline.
Despite this performance, both DistilHuBERT and LightHuBERT consume significantly more MACs per second of speech due to retaining the convolutional layers from the original HuBERT model. Specifically, in LightHuBERT, convolutional layers account for approximately 64% of MACs while contributing only 4.4% of the model’s total parameters. In contrast, the convolutional layers of FitHuBERT, ARMHuBERT, and STaRHuBERT are optimized, leading to significantly lower MACs per second.
@@ V-C Contextualizing prior work
While these five methods have a similar number of parameters, their phone recognition performance varies significantly. The difference is primarily attributed to model depth—DistilHuBERT has only two layers, whereas FitHuBERT, LightHuBERT, ARMHuBERT, and STaRHuBERT have twelve. This result indicates that for phone recognition, the depth of Transformers is more critical than the number of heads or the size of the feed-forward layers.
However, in terms of actual runtime, LightHuBERT is the slowest, while FitHuBERT offers no significant improvement over simply using the first six layers of MelHuBERT. As discussed in the previous section, increasing the number of layers results in more GPU kernel calls, where the overhead of these calls can become a bottleneck, especially when each call is relatively lightweight. Consequently, DistilHuBERT achieves a notable runtime reduction by limiting the model to two layers.
ARMHuBERT demonstrates a clear runtime advantage over STaRHuBERT, despite their similar MACs and parameter counts. This advantage arises from ARMHuBERT’s ability to reuse attention maps for every two layers, significantly reducing the computation cost of self-attention—a result that aligns with our analysis in the previous section.
In summary, our analysis contextualizes the performance of DistilHuBERT, FitHuBERT, LightHuBERT, ARMHuBERT, and STaRHuBERT, providing insights into their respective strengths and weaknesses.
V-D Combining compression techniques
@@
VI Conclusion
Based on the weight pruning results, we can conclude that there exists small but strong self-supervised models that can produce representations that generalize across tasks. Weight pruning, even the vanilla version used in this work, is in fact able to find models much smaller than the recent approaches. However, smaller models do not necessarily run faster. Even though runtime is correlated with MACs, the characteristics of the hardware (such as the amount of parallelization) plays a much important role. In other words, if the goal is to improve runtime, reducing the number of GPU kernel calls is much more effective than reducing the size of the kernel calls. We hope that this study encourages future work to consider multiple dimensions when designing and evaluating compression strategies.